OpenStudy (anonymous):

y=log8(48x^3-2x^5)+log8(x) . 8 is the base. Solve and find the sum of the natural number x values.

jhonyy9 (jhonyy9):

us log8(a) +log8(b)=log8(a*b)

OpenStudy (anonymous):

And after that?

jhonyy9 (jhonyy9):

so than will get y=log8(x(48x^3 -2x^5)) =log8(48x^4 -2x^6)= =log8(2x^4(24-x^2))

OpenStudy (asnaseer):

there must be some information missing - this cannot be "solved". e.g. is the question solve for x when y=0?

OpenStudy (anonymous):

The problem was the same I wrote, but I guess it's supposed to be y=0.

OpenStudy (experimentx):

i thought the same. but says to " find the sum "

OpenStudy (anonymous):

Answer is supposed to be 10.

OpenStudy (experimentx):

use this property, log8 x = y => x = 8^y = 1 ... rest is quite easy.

OpenStudy (anonymous):

I'm kind of stuck there, don't remember the way to solve that equation.

OpenStudy (anonymous):

2x^6-48x^4+1=0

OpenStudy (asnaseer):

the question needs to be checked otherwise you could spend time solving the "wrong" question.

OpenStudy (anonymous):

y=log8(48x^3-2x^5)+log8(x) . 8 is the base.Find the argument values. And then the sum of natural number solutions.

OpenStudy (experimentx):

didn't you say ... y=0??

OpenStudy (anonymous):

The problem given is in the form of y=log8(48x^3-2x^5)+log8(x). I suppose y might be 0?

jhonyy9 (jhonyy9):

ErkoT - so i think you need to check the interval where the logarithm function is defined firstly and after us this interval for this logarithm with base 8 what you have got in the last step

OpenStudy (experimentx):

lol, you will never be able to solve it. that's a equation in x and y, that will give you are curve.

OpenStudy (anonymous):

And now I'm officially confused.

jhonyy9 (jhonyy9):

???

OpenStudy (anonymous):

Never mind with that problem, I'll put another one to the left.

jhonyy9 (jhonyy9):

first you need to say where one logarithm function is defined ? can you tel me now here ?

OpenStudy (anonymous):

I think I understand that part, but then I have no idea.

jhonyy9 (jhonyy9):

why ? how you have learned on your math study in your school ?

OpenStudy (anonymous):

I'm just tired as hell and english is not my native language so it's pretty difficult to translate all of the text in both ways.

jhonyy9 (jhonyy9):

so if i remember correct need to be allways greater than zero ...

OpenStudy (asnaseer):

ok - I think jhonyy9 has interpreted the question correctly. you need to find the sum of the "integer" solutions to this equation. as jhonyy9 indicated, you first need to determine the valid range.

OpenStudy (anonymous):

Yes, makes sense.

OpenStudy (asnaseer):

it's a very badly worded question though - kudos to jhonyy9 for getting it right!

jhonyy9 (jhonyy9):

camon asnaseer hope you will can make it right till the end

OpenStudy (anonymous):

Sorry for that, not that good at translating from one language to another.

OpenStudy (asnaseer):

I would suggest you also post in original language to avoid confusion like this in the future (as others may be able to translate it better)

OpenStudy (anonymous):

Note taken.

OpenStudy (asnaseer):

so we need to find the sum of the integer solutions to this:\[log_8(2x^4(24-x^2))>0\]

OpenStudy (asnaseer):

which can be simplified to:\[2x^4(24-x^2)\gt8^0\]or,\[2x^4(24-x^2)\gt1\]

OpenStudy (anonymous):

And how's that one solved?

OpenStudy (anonymous):

I'm sorry if the questions are too stupid, but I'm just overly tired and still a lot to do.

OpenStudy (asnaseer):

I think we can use this approach:\[2x^4(24-x^2)-1\gt0\]then use difference of two squares to get:\[(x^2\sqrt{2(24-x^2)}+1)(x^2\sqrt{2(24-x^2)}-1)\gt0\]

jhonyy9 (jhonyy9):

asnaseer - what sign ,,kudos" ?

OpenStudy (asnaseer):

jhonyy9: see here: http://en.wikipedia.org/wiki/Kudos

jhonyy9 (jhonyy9):

ok thank you

OpenStudy (asnaseer):

yw

jhonyy9 (jhonyy9):

nice very

OpenStudy (anonymous):

So I get 48x^4-48x^6-1=0 , how to solve that one?

jhonyy9 (jhonyy9):

check it newly because what you have got there not is right

jhonyy9 (jhonyy9):

yes exactly

jhonyy9 (jhonyy9):

ErkoT can you solve one inegality ?

OpenStudy (anonymous):

Yes, got the same, stupid mistake.

OpenStudy (anonymous):

And how do you solve that one?

OpenStudy (asnaseer):

ok, so we want to solve this inequality:\[48x^4−2x^6−1\gt0\]

OpenStudy (anonymous):

Yes.

OpenStudy (asnaseer):

and it asked for "natural" numbers - so I am guessing this implies we only want the positive x values

OpenStudy (anonymous):

I agree.

OpenStudy (asnaseer):

ok - one way I can think of is to find the maximum value of this function and then the solutions are just all natural numbers less than or equal to this value

OpenStudy (anonymous):

Derive it?

OpenStudy (anonymous):

Or differentiate or I don't even know the exact word.

OpenStudy (asnaseer):

yes, differentiate to get:\[-12x^3(x^2-16)=0\]which gives us:\[x=0\]or\[x=\pm4\]

OpenStudy (asnaseer):

we need to differentiate again to find which x gives the maximum.

OpenStudy (asnaseer):

that gives:\[576x^2-60x^4\]

OpenStudy (asnaseer):

which means max is at \(x=\pm4\)so our solution must be: x=1, 2, 3, 4 so sum = 1+2+3+4 = 10

OpenStudy (anonymous):

I think I can't thank you enough for that one! Thanks!

OpenStudy (asnaseer):

you are more than welcome - it was a very interesting problem. I'm glad I could help and thanks to @jhonyy9 for clarifying the question in the first place.

OpenStudy (anonymous):

@jhonyy9 , thank you! Thank you both!

jhonyy9 (jhonyy9):

yw asnaseer was my pleasure good luck bye